Direct Materials, Direct Labor, and Factory Overhead Cost Variance Analysis Mackinaw Inc. processes a base chemical into plastic. Standard costs and actual costs for direct materials, direct labor, and factory overhead incurred for the manufacture of 5,600 units of product were as follows:

Standard Costs Actual Costs

Direct materials 7,300 lb. at $5.30 7,200 lb. at $5.20

Direct labor 1,400 hrs. at $17.00 1,430 hrs. at $17.30

Factory overhead Rates per direct labor hr.,

based on 100% of normal capacity

of 1,460 direct labor hrs.:

Variable cost, $4.50 $6,240 variable cost

Fixed cost, $7.10 $10,366 fixed cost

Each unit requires 0.25 hour of direct labor.

Required:

a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct materials price variance $

Direct materials quantity variance

Total direct materials cost variance $

b. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Direct labor rate variance $

Direct labor time variance

Total direct labor cost variance $

c. Determine variable factory overhead controllable variance, the fixed factory overhead volume variance, and total factory overhead cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Variable factory overhead controllable variance $

Fixed factory overhead volume variance

Total factory overhead cost variance $
Answer #1

Solution a:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
7200 $5.20 $37,440.00 7200 $5.30 $38,160.00 7300 $5.30 $38,690.00
-$720.00 F -$530.00 F
Direct Material Price Variance Direct Material Qty variance
Direct material price variance -$720.00 F
Direct material quantity variance -$530.00 F
Direct material cost variance -$1,250.00 F

Solution b:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
1430 $17.30 $24,739.00 1430 $17.00 $24,310.00 1400 $17.00 $23,800.00
$429.00 U $510.00 U
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $429.00 U
Direct Labor Efficiency variance $510.00 U
Direct labor cost variance $939.00 U

Solution c:

Variable factory overhead controllable variance = Actual cost of variable overhead - Standard cost of variable overhead

= $6,240 - (1400*$4.50) = - $60F

Fixed factory overhead volume variance = Budgeted fixed overhead - Fixed overhead applied

= $10,366 - (1400 * $7.10) = $426 U

Total factory overhead cost variance = - $60 F + $426 U = $366 U